# Benchmarking Retrieval-Augmented Multimomal Generation for Document Question Answering

### v1

@@ 4.1 Evaluation Metric
Method Metric Tokens Quote Selection Multimodal Answer Quality
Prec Rec F1 Prec Rec F1 geL ency Qlty. Coher. Logic uality
Qwen2.5-VL-7B-Inst 7.1k 128 58.0 14.5 23.2 31.3 11.0 16.3 16.6 0.069 0.273 4.05 1.75 1.89 2.36 2.29 2.47
4.2 Baseline Models
Quotes Retrieval.
Multimodal Answer Generation.
We evaluate 60 latest models by using quotes as: (i) multimodal inputs for VLM, and (ii) pure-text inputs for VLM and LLM (see Appendix C.1 for implementation details). Then, we evaluate 5 finetuned models (Qwen2.5 LLMs [59] with 3-72B parameters) using MMDocRAG dev-set. See Appendix C.2 for finetuning details).
4.3 Main Results
Figure 4: Performance difference: base/finetuned models.
@@ 4.3 Main Results
Multimodal:MM Pure-Text:PT MM PT $\Delta\%$ MM PT $\Delta\%$ MM PT $\Delta\%$
Use the same VLM to process both multimodal and pure-text inputs.
Gemini-1.5-Pro 3.8k 3.6k -5.3 59.3 56.2 -5.2 2.77 3.03 +9.4
Gemini-2.0-Pro 3.8k 3.6k -5.3 62.0 62.8 +1.3 3.31 3.51 +6.0
Gemini-2.0-Flash 3.8k 3.6k -5.3 60.0 54.4 -9.3 3.11 3.19 +2.6
Gemini-2.0-Flash-Think 3.8k 3.6k -5.3 66.2 61.0 -7.9 3.61 3.47 -3.9
Gemini-2.5-Pro 3.7k 3.6k -2.7 65.4 65.1 -0.5 3.88 3.79 -2.3
Gemini-2.5-Flash 3.7k 3.6k -2.7 62.4 59.5 -4.6 3.76 3.55 -5.6
Claude-3.5-Sonnet 7.8k 3.8k -51.3 62.5 57.4 -8.2 3.65 3.60 -1.4
GPT-4o-mini 8.5k 3.4k -60.0 56.3 56.6 +0.5 3.60 3.70 +2.8
GPT-4o 6.4k 3.4k -46.9 62.6 57.2 -8.6 3.74 3.69 -1.3
GPT-4.1-nano 14.2k 3.4k -76.1 29.0 40.8 +40.7 3.12 3.39 +8.7
GPT-4.1-mini 9.8k 3.4k -65.3 61.2 61.0 -0.3 3.92 3.90 -0.5
GPT-4.1 6.6k 3.4k -48.5 70.2 68.3 -2.7 4.14 4.07 -1.7
InternVL3-8B 17.1k 3.6k -78.9 37.0 48.1 +30.0 3.14 3.19 +1.6
InternVL3-9B 17.2k 4.0k -76.7 50.9 45.4 -10.8 3.12 3.30 +5.8
InternVL3-14B 17.1k 3.6k -78.9 49.9 49.9 +0.0 3.29 3.48 +5.8
InternVL3-38B 17.1k 3.6k -78.9 53.9 55.0 +2.0 3.45 3.61 +4.6
InternVL3-78B 17.1k 3.6k -78.9 59.8 56.4 -5.7 3.42 3.56 +4.1
Llama4-Scout-17Bx16E 11.6k 3.3k -71.6 38.9 48.2 +23.9 3.13 3.12 -0.3
Llama4-Mave-17Bx128E 11.6k 3.3k -71.6 58.6 58.3 -0.5 3.67 3.59 -2.2
Qw-VL-Plus Qw-Plus 7.1k 3.6k -49.3 25.2 55.4 +119.8 2.36 3.63 +53.8
Qw-VL-Max Qw-Max 7.1k 3.6k -49.3 46.8 58.9 +25.9 3.35 3.77 +12.5
QVQ-Max QwQ-Plus 6.8k 3.6k -47.1 12.3 59.6 +384.6 3.36 3.63 +8.0
Qw2.5-VL-7B Qw2.5-7B 7.1k 3.6k -49.3 16.6 45.8 +175.9 2.47 3.34 +35.2
Qw2.5-VL-32B Qw2.5-32B 7.0k 3.6k -48.6 36.2 58.9 +62.7 3.73 3.63 -2.7
Qw2.5-VL-72B Qw2.5-72B 7.1k 3.6k -49.3 57.5 59.1 +2.8 3.47 3.75 +8.1
Table 5: Using 20 quotes for multimodal generation. $\Delta\%$ is calculated by values (PT-MM)/MM in percentage.
4.4 Multimodal vs Pure-text Quotes: Comparison and Analysis
As shown in Table 5 and Table 8, we compare model performance when quotes are provided as either pure-text or multimodal inputs. Multimodal quotes significantly increase token usage, as images are typically encoded with more tokens. Interestingly, Gemini models maintain similar token usage across both modes, indicating efficient image encoding. Gemini, Claude, and GPT models demonstrate superior quote selection performance in the multimodal setting and comparable answer quality across both input types. In contrast, Qwen models perform significantly better in both quote selection and answer generation when using pure-text inputs. Smaller VLMs, compared to their LLM counterparts, struggle to effectively process long multimodal input sequences. For instance, the Qwen-7B and 32B LLMs achieve 175.9% and 62.7% higher F1 scores for quote selection, respectively, compared to their equivalent VLMs. Further qualitative analysis is provided in Appendix F.2.
Method Metric Image Quote F1 Answer Avg.
VLM OCR $\Delta$ VLM OCR $\Delta$
15 Quotes Qwen2.5-7B-Inst 59.8 49.6 -10.2 3.37 3.15 -0.22
Llama3.1-8B-Inst 61.1 52.4 -8.7 3.33 3.19 -0.14
Llama3.3-70B-Inst 71.8 64.1 -7.7 3.14 3.08 -0.06
Qwen2.5-72B-Inst 73.3 65.7 -7.6 3.76 3.55 -0.21
Qwen-Max 74.4 65.7 -8.7 3.77 3.63 -0.14
Deepseek-V3 76.5 70.2 -6.3 3.75 3.68 -0.07
Gemini-2.0-Pro 77.4 74.9 -2.5 3.50 3.41 -0.09
Gemini-2.0-Fl-Tk 74.9 73.0 -1.9 3.51 3.46 -0.05
GPT-4o 71.6 69.4 -5.9 3.73 3.65 -0.08
Avg. results 71.7 65.0 -6.7 3.54 3.42 -0.12
20 Quotes Qwen2.5-7B-Inst 53.5 43.5 -10.0 3.34 3.15 -0.19
Llama3.1-8B-Inst 52.2 45.8 -6.4 3.25 3.16 -0.09
Llama3.3-70B-Inst 65.1 60.6 -4.5 3.24 3.13 -0.11
Qwen2.5-72B-Inst 68.0 59.7 -8.3 3.75 3.50 -0.25
Qwen-Max 69.3 59.9 -9.4 3.77 3.62 -0.15
Deepseek-V3 71.8 65.8 -6.0 3.74 3.59 -0.15
Gemini-2.0-Pro 77.0 70.9 -6.1 3.51 3.32 -0.18
Gemini-2.0-Fl-Tk 73.5 68.3 -5.2 3.47 3.41 -0.06
GPT-4o 66.4 63.8 -2.6 3.69 3.59 -0.10
Avg. results 66.3 59.8 -6.5 3.53 3.39 -0.14
4.5 Multimodal Quotes as text: VLM-text vs OCR-text
4.6 Quotes Selection Analysis
Figure 6: Quotes selection accuracy at all positions.
Method Recall@10 Recall@15 Recall@20
Txt Img Txt Img Txt Img
Text DPR 25.5 53.5 31.4 59.9 35.9 63.9
ColBERT 37.4 64.1 42.8 69.1 46.0 72.8
BGE 38.8 64.9 43.6 70.2 47.0 74.2
E5 41.7 63.5 46.4 69.1 49.5 73.7
Contriever 37.7 64.1 42.8 69.8 46.8 73.4
GTE 38.2 63.0 43.3 69.1 47.3 72.8
Visual DSEwiki-ss 24.7 67.1 29.6 75.3 33.4 79.9
DSEdocmatix 25.6 65.7 30.1 75.0 33.7 78.2
ColPali 27.3 68.2 32.6 77.5 35.2 81.2
ColQwen 28.5 70.8 33.7 79.2 36.0 84.3
Hybrid ColP+ColB 38.2 67.3 42.6 79.2 46.8 83.4
ColP+BGE 39.2 67.3 43.5 79.2 47.7 83.4
ColQ+ColB 38.2 68.5 42.6 81.0 46.8 85.2
ColQ+BGE 39.2 68.5 43.5 81.0 47.7 85.2
Table 7: Retrieval Results.
4.7 Quotes Retrieval Results
5 Related Work
Interleaved Text-Image Generation
Multimodal RAG and Benchmarks.
Retrieval-Augmented Generation (RAG) retrieves relevant quotations as context for answer generation [34]. Multimodal RAG (MRAG) extends RAG by retrieving and leveraging multimodal knowledge (e.g., image-text pairs) for VQA [5, 37]. MuRAR [83] tackles source attribution by retrieving multimodal elements from webpage. M2RAG [42] builds upon MuRAR by proposing a multi-stage image insertion framework that uses model multiple times during answer generation. Although MuRAR and M2RAG enable multimodal answer generation, their benchmarks are limited to webpage domain and lack annotations of supporting evidence.
DocVQA and DocRAG Benchmarks.
Early DocVQA benchmarks focus on single-page VQA, such as DocVQA [43], InfoVQA [44], and TAT-DQA [81]. To mitigate the limitation of single-page input, DUDE [33], MP-DocVQA [70], SildeVQA [66] extend context lengths to averages 5.7, 8.3, and 20 pages respectively. Two most recent MMLongBench-Doc [41] and DocBench [84], formulate DocVQA as long-context tasks by inputting entire documents (averaging 50-70 pages). To address increasing document length, M3DocVQA [10], M-Longdoc [9], and MMDocIR [16] propose DocRAG tasks, incorporating evidence retrieval followed by answer generation over the retrieved multimodal evidence. To the best of our knowledge, no existing DocVQA or DocRAG benchmarks focus on multimodal interleaved generation.
6 Conclusion
Appendix Overview
The appendix includes the following sections:
  • Appendix B: Provides supplementary experimental results, with (B.1) related results by using 15 quotes for multimodal generation, (B.2) comprehensive results comparing thinking and non-thinking modes, (B.3) comprehensive results comparing different models using OCR and LLM text for image quote representation, and (B.4) fine-grained results by document type for quote selection and multimodal answer generation.
  • Appendix H: Discusses the limitations of MMDocRAG.
Appendix A Evaluation Metric of Multimodal Answer Generation
A.1 Related Work of multimodal generation
Multimodal generation, particularly interleaved image-text sequence generation, involves generating outputs that integrate visual and textual information in a cohesive manner (see Section 5). This capability facilitate applications such as storytelling, question answering, and document comprehension. Recent benchmark, MM-Interleaved [69], MMIE [75], GATE Opening [80], and M2RAG [42] provide comprehensive evaluations for multimodal generation. Commonly adopted metrics include fluency, relevance, image-text coherence, and content quality. These are evaluated through human annotation or automated scoring using large language models such as GPT-4. Specifically, fluency assesses the grammatical correctness and readability of text, relevance measures the alignment of generated content with the prompt, image-text coherence evaluates the logical connection between images and text, and content quality addresses the completeness and richness of the output. Our benchmark, MMDocRAG, adopts established metrics such as fluency, image-text coherence, and content quality. Additionally, we incorporate BLEU [53] and ROUGE-L [36] scores to quantitatively assess the semantic similarity between generated and gold answers.
A.2 Evaluation Metrics: details and implementations
1. Quote Selection Metrics.
2. Surface-level Similarity Metrics.
BLEU $=BP\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right),$ where $\text{where }BP$ =$=\begin{cases}1,&\text{if }c>r\\ \exp\left(1-\frac{r}{c}\right),&\text{if }c\leq r\end{cases}$ (2)
[…]
where $R_{LCS}$ and $P_{LCS}$ are the recall and precision based on LCS length, and ||$\cdot$|| ⋅ | refers to the length of generated or reference answer. $\beta$ is typically set to favor recall ($\beta=1.2$ by default).
3. LLM-as-Judge Evaluation Criteria.
Method In-token Usage Overall Quote F1 Answer Avg.
Multimodal (MM) Pure-Text (PT) MM PT $\Delta\%$ MM PT $\Delta\%$ MM PT $\Delta\%$
@@ A.2 Evaluation Metrics: details and implementations
Table 8: Using 15 quotes for multimodal generation. $\Delta\%$ is calculated by values (PT-MM)/MM and displayed in percentage.
@@ B.4 Fine-grained Results by Document Domains
As illustrated in Figure 7, different models exhibit distinct performance patterns across various document types. Our findings include:
@@ C.2 Implementation Details of LLM Finetuning
As described in Section 4.2, we finetune five Qwen2.5 LLMs: Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct, and Qwen2.5-72B-Instruct.
@@
Supervised fine-tuning is performed with the SWIFT framework, utilizing memory-efficient methods such as LoRA [26], FlashAttention [11], and DeepSpeed [60]. We set the LoRA rank to 16 and alpha to 32. Given the average input length of 3.6k tokens (see Table 4), the maximum sequence length is set to 8k. Training is performed for one epoch, using gradient accumulation to update LoRA weights every 8 training steps. Our parameter-efficient finetuning can be done using 4 A100 GPUs.
@@ C.3 Implementation Details of Retriever
{tabu}
l|l|c|c|l Model Dimension Base Model HuggingFace Checkpoint
Text
DPR [31] 768 BERT-base [14] facebook/dpr-ctx_encoder-multiset-base facebook/dpr-question_encoder-multiset-base
@@
Visual
DSEwiki-ss [40] 3,072 Phi-3-Vision [1] Tevatron/dse-phi3-v1.0
@@
Text retrieval methods are typically categorized into sparse and dense retrieval. Sparse retrievers, such as TF-IDF [63] and BM25 [62], compute relevance based on word frequency statistics, with BM25 adding nonlinear frequency saturation and length normalization. Dense retrievers represent content as vectors: DPR [31] is a pioneering work for QA tasks; ColBERT [32] enables efficient late interaction for fine-grained question-document matching; Contriever [28] employs contrastive learning to enhance dense representations; E5 [72] and BGE [77] introduce improved training and data strategies; and GTE [35] incorporates graph-based methods for further enhancement. Despite recent progress, most text retrievers overlook valuable visual information that may be embedded in documents.
@@
In our experiments (section 4), we implement 6 dense text retrievers: DPR [31], ColBERT [32], Contriever [28], E5 [72], BGE [77], and GTE [35]. All models use the BERT WordPiece tokenizer and a maximum sequence length of 512 tokens [14]. We utilize publicly available checkpoints from HuggingFace (see Table C.3 for details) and the sentence-transformers library [^7] https://www.sbert.net/ for deploying E5, BGE, and GTE.
@@
We implement four visual retrievers: DSEwiki-ss [40], DSEdocmatix [40], ColPali [17], and ColQwen [17]. These models use image tokenizers to convert image quotes into $14\times 14$ pixel patches, each corresponding to a visual token. We employ pre-trained checkpoints from HuggingFace, with configuration details listed in Table C.3.
@@ 3.1 Answer Output Format:
- Entries that describe situations identical to those in the long answer but lack critical keys should also not be selected; - If there is no relevant text passages, set "need_text"=False; - If there are any relevant text passages, set "need_text"=True; - If "need_text"==True, please select the relevant text by choosing the text passage indices; - Note: Do not forcibly select evidence, only select evidence that is fully or strongly relevant. If there is no such evidence, then it should be considered as having no evidence, avoid making forced associations just to select evidence; # 3. Citing/Quoting Text Passage Indices in Long Answer
@@ Appendix G License Agreement
Appendix H Limitations
The limitations of our methods are as follows:
  • Limited Multimodal Retrieval Capability: MMDocRAG primarily focuses on multimodal answer generation by evaluating the ability of LLMs and VLMs to select and integrate relevant multimodal quotes from a noisy set. In our proposed MMDocRAG, each question is associated with an average of 2.7 gold quotes out of 15/20 candidates, resulting in only 18.0/13.5% relevant quotes for answer generation. However, real-world RAG scenarios might present cases where no relevant quotes are retrieved, especially with imperfect multimodal retrieval modules. Our current setup does not adequately capture such rare unanswerable cases.
  • Fine-tuning Limited to LLMs: Our experiments fine-tune five Qwen2.5 models (ranging from 3B to 72B parameters) where multimodal quotes are linearized as pure text sequences. We do not report fine-tuning results on Qwen2.5-VL models due to computational constraints: multimodal input sequences are significantly longer (7.1k vs.3.6k tokens on average), and, in extreme cases with high-resolution images, can exceed 20k tokens. While we attempted to fine-tune Qwen2.5-VL models, training was unstable and frequently interrupted. As a result, we do not present VLM fine-tuning results in this work.

### v2

@@ A.2 Evaluation Metrics: details and implementations
Multimodal (MM) Pure-Text (PT) $M$ PT $\Delta\%$ $M$ PT $\Delta\%$ $M$ PT $\Delta\%$
@@
Table 9: Using 15 quotes for multimodal generation. $\Delta\%$ is calculated by values (PT-MM)/$M$ and displayed in percentage.
@@ B.4 Fine-grained Results by Document Domains
Beyond main results (Section 4.3) on MMDocRAG, we present fine-grained results breakdown by domains. As illustrated in Figure 8, different models exhibit distinct performance patterns across various document types. Our findings include:
@@
B.5 Fine-grained Results by Question Types
Beyond main results (Section 4.3) on MMDocRAG, we present fine-grained results breakdown by question types. As illustrated in Figure 9, different models exhibit distinct performance patterns across various question types. Our findings include:
  • All models achieve highest performance in "Descriptive" and "Comparative" categories, attributed to their straightforward information extraction requirements. Models perform worst on "Interpretative" and "Inferential" categories due to increased reasoning complexity.
  • Advanced VLMs (GPT-4.1, Gemini2.5-pro) consistently achieve higher and more balanced scores across question types, especially in complex reasoning categories. This indicates superior multi-step reasoning capacity compared to smaller models that show pronounced degradation with increased question complexity.
  • Answer quality positively correlates with F1 score of quote selection across all question types, with strongest correlation in "Analytical" and "Comparative" categories. F1 score reflects evidence identification ability while answer quality measures generation capability from selected evidence.
  • GPT-4.1 exhibits the most balanced performance across both metrics, maintaining high scores even for complex questions. Gemini2.5-pro excels in "Descriptive" tasks, Claude-3.5-sonnet shows challenges in complex reasoning, and Llama4-Mave-17Bx128E displays the most constrained performance envelope across all question types.
B.6 Fine-grained Results by Evidence Types
Beyond main results (Section 4.3) on MMDocRAG, we present fine-grained results breakdown by evidence types. Figure 10 reveals how different evidence configurations impact model performance in multimodal RAG tasks. Our analysis yields several key findings:
  • Single vs. Multiple Image Evidence: All models consistently achieve higher F1 scores and answer quality when questions require evidence from a single image rather than multiple images. This pattern indicates that synthesizing information across multiple visual sources presents a significant challenge for current VLMs.
  • Single vs. Multiple Page Evidence: Questions with evidence contained within a single page consistently outperform those requiring multi-page evidence across all models. This suggests that information gathering and consolidation across document boundaries remains a substantial bottleneck.
  • Single vs. Cross-Modal Evidence: Unlike the previous patterns, cross-modal evidence preferences vary by model architecture. GPT-4.1 and Llama4-17Bx128 perform better with single-modal evidence, while Gemini2.5-pro and Claude-3.5-sonnet show superior performance with cross-modal evidence. This divergence reflects fundamental differences in how these models handle modality fusion and integration.
  • Overall Model Performance: GPT-4.1 maintains the highest performance across all evidence configurations, demonstrating robust scalability as evidence complexity increases. Gemini2.5-pro shows particularly strong gains in cross-modal settings, while Claude-3.5-sonnet and Llama4-17Bx128 exhibit more constrained performance envelopes, with Llama4 showing the most limited adaptability to evidence complexity variations.
@@ B.6 Fine-grained Results by Evidence Types
Figure 9: The fine-grained (by question types) results of 4 representative large models in two settings: using 20 quotes as either pure-text or interleaved manner. We show the F1 score of quotes selection (ranging from 50 to 75) and answer quality (ranging from 3.4 to 4.4).
(a) Performance comparison between questions with single and multiple image quotes as evidence.
(b) Performance comparison between questions with single and multiple pages as evidence.
(c) Performance comparison between questions with single and cross modal evidence type.
Figure 10: Fine-grained results of 4 VLMs using both 20 multimodal ($M$) and pure-text (PT) quotes. The breakdown is according to questions consisting of: (a) single or multiple image quotes as evidence, (b) single or multiple pages as evidence, and (c) single or cross modal evidence type.
@@ C.2 Implementation Details of LLM Finetuning
As described in Section 4.2, we finetune (i) five Qwen2.5 LLMs including: Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct, and Qwen2.5-72B-Instruct, (ii) two Qwen2.5-VL VLMs namely: Qwen2.5-VL-3B-Instruct and Qwen2.5-VL-7B-Instruct, and (iii) two InternVL-3 VLMs namely: InternVL-3-8B and InternVL-3-9B.
@@
Supervised fine-tuning is performed with the SWIFT framework, utilizing memory-efficient methods such as LoRA [31], FlashAttention [11], and DeepSpeed [66]. We set the LoRA rank to 16 and alpha to 32. For finetuning LLMs, we set the maximum sequence length to 8k, given the average input length of 3.6k tokens (see Table 3). For finetuning VLMs, we set the maxium sequence length to 32k instead, given that images need more tokens for accurate representation. Training is performed for one epoch, using gradient accumulation to update LoRA weights every 8 training steps.
@@ C.3 Implementation Details of Retriever
Model Dimension Base Model HuggingFace Checkpoint
Text DPR [36] 768 BERT-base [14]
facebook/dpr-ctx_encoder-multiset-base
facebook/dpr-question_encoder-multiset-base
@@
Visual DSEwiki-ss [46] 3,072 Phi-3-Vision [1] Tevatron/dse-phi3-v1.0
@@
Text retrieval methods are typically categorized into sparse and dense retrieval. Sparse retrievers, such as TF-IDF [69] and BM25 [68], compute relevance based on word frequency statistics, with BM25 adding nonlinear frequency saturation and length normalization. Dense retrievers represent content as vectors: DPR [36] is a pioneering work for QA tasks; ColBERT [37] enables efficient late interaction for fine-grained question-document matching; Contriever [33] employs contrastive learning to enhance dense representations; E5 [78] and BGE [83] introduce improved training and data strategies; and GTE [41] incorporates graph-based methods for further enhancement. Despite recent progress, most text retrievers focus on textual content [40] and overlook valuable visual information that may be embedded in documents.
@@
In our experiments (section 4), we implement 6 dense text retrievers: DPR [36], ColBERT [37], Contriever [33], E5 [78], BGE [83], and GTE [41]. All models use the BERT WordPiece tokenizer and a maximum sequence length of 512 tokens [14]. We utilize publicly available checkpoints from HuggingFace (see Table 14 for details) and the sentence-transformers library [^7] https://www.sbert.net/ for deploying E5, BGE, and GTE.
@@
We implement four visual retrievers: DSEwiki-ss [46], DSEdocmatix [46], ColPali [21], and ColQwen [21]. These models use image tokenizers to convert image quotes into $14\times 14$ pixel patches, each corresponding to a visual token. We employ pre-trained checkpoints from HuggingFace, with configuration details listed in Table 14.
@@ 3.1 Answer Output Format:
     - Entries that describe situations identical to those in the long answer but lack critical keys should also not be selected; - If there is no relevant text passages, set "need_text"=False; - If there are any relevant text passages, set "need_text"=True; - If "need_text"==True, please select the relevant text by choosing the text passage indices; - Note: Do not forcibly select evidence, only select evidence that is fully or strongly relevant. If there is no such evidence, then it should be considered as having no evidence, avoid making forced associations just to select evidence; # 3. Citing/Quoting Text Passage Indices in Long Answer